# TalkingHeadBench: A Multi-Modal Benchmark & Analysis of Talking-Head DeepFake Detection

### v1

@@ Abstract
The rapid advancement of talking-head deepfake generation fueled by advanced generative models has elevated the realism of synthetic videos to a level that poses substantial risks in domains such as media, politics, and finance. However, current benchmarks for deepfake talking-head detection fail to reflect this progress, relying on outdated generators and offering limited insight into model robustness and generalization. We introduce TalkingHeadBench, a comprehensive multi-model multi-generator benchmark and curated dataset designed to evaluate the performance of state-of-the-art detectors on the most advanced generators. Our dataset includes deepfakes synthesized by leading academic and commercial models and features carefully constructed protocols to assess generalization under distribution shifts in identity and generator characteristics. We benchmark a diverse set of existing detection methods, including CNNs, vision transformers, and temporal models, and analyze their robustness and generalization capabilities. In addition, we provide error analysis using Grad-CAM visualizations to expose common failure modes and detector biases. TalkingHeadBench is hosted on https://huggingface.co/datasets/luchaoqi/TalkingHeadBench with open access to all data splits and protocols. Our benchmark aims to accelerate research towards more robust and generalizable detection models in the face of rapidly evolving generative techniques.
@@ 1 Introduction
Deepfake technology, known for its ability to manipulate facial features and produce highly realistic synthetic videos, has seen rapid adoption in entertainment liu_deepfacelab_2023 ; abdolahnejad_deep_2020 , marketing kaur_deepfake_2024 , and personalized content creation qi_mytimemachine_2024 ; qi_my3dgen_2025 ; wei_personalized_2025 . Yet, these advancements also introduce serious societal risks. Among deepfake types, talking-head deepfakes, which generate highly realistic head-and-shoulder videos of individuals speaking, are especially concerning. Driven by advanced generative models, these videos are potent tools for spreading misinformation, manipulating public discourse, and executing social engineering attacks gambin_deepfakes_2024 ; saini_deepfake_2024 ; saif_fake_2024 . Their realism and ease of distribution make them particularly dangerous in sensitive domains like politics, finance, and media. In one real-world case, a finance worker was tricked into transferring \$25 million during a video call with a deepfake impersonating the company’s CFO magramo_finance_2024 . These growing threats highlight the urgent need to understand talking-head deepfakes and to develop robust detection and mitigation strategies gambin_deepfakes_2024 ; sirra_csslno_2024 ; liu_evolving_2025 .
@@
In response to the growing threats posed by talking-head deepfakes, numerous facial deepfake benchmarks have been proposed (see Tab. 1, which played a central role in advancing detection methods. However, most existing benchmarks either focus on face-swapping techniques that are easy to detect rossler2019faceforensics++ , or rely on uncurated ‘in-the-wild’ datasets chandra2025deepfake ; yan2024df40 that include older, GAN-based deepfake rossler2019faceforensics++ or lip-sync-specific generators hou2024polyglotfakenovelmultilingualmultimodal ; cai2024av ; cai2022you . These datasets often lack diversity in generation methods, suffer from low data quality due to artifacts, and do not provide controlled multi-modal evaluation settings. Furthermore, they do not reflect the capabilities of recent diffusion-based talking-head generators, which produce more realistic manipulations by synthesizing the entire facial region and controlling both pose and expression. To address these gaps, we introduce TalkingHeadBench—the first manually curated, multi-generator, multi-modal benchmark designed specifically for evaluating detector robustness against state-of-the-art talking-head deepfakes.
@@
Our benchmark features high-quality synthetic videos generated using six academic and one commercial talking-head generator using both audio and video-driven diffusion and transformer-based models  drobyshev2024emoportraits ; cui2024hallo2 ; xu2024hallo ; wei2024aniportrait ; guo2024liveportrait . We also carefully design evaluation protocols to assess the generalization and robustness of state-of-the-art (SOTA) deepfake detectors dong2023implicit ; shao2025deepfake ; xu2023tall ; xu2024learning ; liu2024lips across identity and generator shifts in training and testing data. In addition to benchmarking, we also present a detailed error analysis using Grad-CAM Selvaraju_2019 to identify common failure cases of SOTA detectors on different generators and inform future improvements in detection strategies. We then note the limitations of performance of SOTA detectors and identify areas of improvement, and provide guidelines on what performance future detectors should aspire to and how to analyze them.
@@
TalkingHeadBench benchmark is currently hosted in the HuggingFace repository, which provides train-test data for all generators across multiple protocols. In the near future, we promise to develop a more active benchmark repository by a) developing the API/code/metrics that also allow authors of new generators to evaluate the ‘detectability’ of their generated videos; b) setting up two leaderboards, which compare the performance of generators and detectors, respectively; c) updating the list of generators and detectors once every 6 months, to accelerate the development cycle.
We summarize our main contributions as follows:
  • We introduce TalkingHeadBench, a multi-modal multi-generator benchmark for talking-head deepfake detection, a significantly more challenging benchmark than previously used FaceForensics++ rossler2019faceforensics++ , where SOTA detectors easily achieve near-perfect accuracy.
  • We design evaluation frameworks to explicitly study detector robustness and generalization across train-test distribution shift caused by varying identity and generator properties. Our analysis indicates that even the best detectors lack robustness, indicating the need for future research in developing advanced detectors using our benchmark.
  • We explicitly analyze why SOTA detectors fail on certain generators using Grad-CAM, providing valuable insights into their biases and presenting directions for future research.
@@
Table 1: Overview of Facial Deepfake Benchmarks. TalkingHeadBench is the first manually curated, multi-modal, multi-generator benchmark focused on talking-head deepfakes, specifically designed to evaluate detector robustness and generalization. Curation removes $\sim$60of low-quality samples with obvious artifacts, ensuring a more realistic and challenging benchmark. Generators marked with ✓* use lip-sync-only techniques.
@@
DeepFake-eval-2024 chandra2025deepfake N/A Video/Audio/Image 2025 1,208 767
FakeAVCeleb khalid2021fakeavceleb Multiple Audio/Video 2021 500 19,500
FaceForensics++ rossler2019faceforensics++ Multiple Video 2019 1,000 4,000
LAV-DF cai2022you ✓* Single Audio/Video 2022 36,431 99,873
AV-Deepfake1M cai2024av ✓* Single Audio/Video 2023 286,721 860,039
PolyGlotFake hou2024polyglotfakenovelmultilingualmultimodal ✓* Multiple Audio 2024 766 14,472
DF-40  yan2024df40 Multiple Video/Image 2024 1,590 1M+
TalkingHeadBench (Ours) Multiple Audio/Video 2025 2,312 2,984
@@ 2 Related Work
Deepfake generation has leveraged various facial manipulation techniques over the years. Among them, face swapping methods nirkin2019fsgan ; rosberg2023facedancer ; chen2020simswap are widely used, combining facial regions of various sources to create synthetic videos. They often introduce visible artifacts such as boundary inconsistency, lighting mismatch, or texture irregularity, which modern detectors can learn to exploit easily.
@@
More recently, talking-head generation has emerged as a more sophisticated approach. These models synthesize the full face, including expressions and lip motion, from a static portrait using audio or video driving signals—eliminating the obvious seam-based flaws of earlier methods. SOTA diffusion-based talking-head models (see Tab. 2) produce highly realistic outputs with synchronized audio-visual features, significantly raising the difficulty for detection algorithms.
@@
Despite these advances, existing deepfake benchmarks remain limited in scope. As shown in Tab. 1, most prior datasets either focus on older GAN-based generators or target narrow manipulation types such as face swapping (e.g., FakeAVCeleb khalid2021fakeavceleb and FaceForensics++ rossler2019faceforensics++ ) or scrape social media content without controlling of generators (e.g., DeepFake-eval-2024 chandra2025deepfake ). While newer datasets like LAV-DF cai2023glitch , AV-Deepfake1M cai2024av , and PolyGlobFake hou2024polyglotfakenovelmultilingualmultimodal incorporate talking-head deepfakes, they mostly emphasize lip-sync-driven synthesis and lack broader modality or generator diversity. Further, these datasets are often massive but uncurated, resulting in many low-quality samples. In contrast, our dataset is built using a curated collection of cutting-edge diffusion-based talking-head generators spanning both audio and video-driven modalities. Through manual inspection, we discarded over 60% of generated videos due to visible artifacts (see Tab. 4), ensuring that only high-quality, realistic samples remain—thereby offering a more meaningful and challenging benchmark.
@@
Recent detection methods (see Tab. 3), though effective on static and unimodal datasets like FaceForensics++ rossler2019faceforensics++ , have not been thoroughly evaluated against multi-modal diffusion-based deepfakes. To fill this gap, we introduce a comprehensive benchmark that reflects the current generation landscape. Our results show that SOTA detectors, which achieve near-perfect accuracy on traditional benchmarks like FaceForensics++, struggle significantly when tested on our dataset, highlighting poor robustness to identity and generator shifts and calling for stronger generalization capabilities.
@@
Table 2: Overview of talking-head generators, highlighted were used for creating TalkingHeadBench.
@@
Image, Audio Diffusion Hallo xu2024hallo arXiv24
Image, Audio Diffusion Hallo2 cui2024hallo2 ICLR25
Image, Audio, Text Diffusion Hallo3 cui2024hallo3 CVPR25
Image, Video Diffusion X-Portrait xie2024x SIGGRAPH24
Image, Video Non-diffusion LivePortrait guo2024liveportrait arXiv24
Image, Video Non-diffusion EMOPortraits drobyshev2024emoportraits CVPR24
Image, Video Non-diffusion MCNet hong2023implicit ICCV23
Image, Video Non-diffusion FSRT rochow2024fsrt CVPR24
Image, Audio/Video Diffusion AniPortrait wei2024aniportrait arXiv24
Image, Audio/Video Diffusion SkyReels-A1 qiu2025skyreelsa1expressiveportraitanimation arXiv25
Image, Text Diffusion MAGI-1 magi1 N/A
@@
Image CNN SBI shiohara2022detecting CVPR22
Image Neural Network CADDM dong2023implicit CVPR23
Image Neural Network HiFiNethifi_net_xiaoguo //HiFiNet++xiao_hifinet_plusplus CVPR23/IJCV24
Image Transformer DeepFake-Adapter shao2025deepfake IJCV25
Image Transformer GenConVit deressa2025genconvitdeepfakevideodetection arXiv25
Image Transformer RFFR shi2025real PR25
Image Neural network LAA-Net nguyen2024laa CVPR24
Video Transformer TALL xu2023tall /TALL++ xu2024learning ICCV23/IJCV24
Video CNN AltFreezing wang2023altfreezing CVPR23
Video CNN/Transformer NACO zhang2024learning ECCV24
Video Neural network MM-Det song2025learningmultimodalforgeryrepresentation NeurIPS24
Video Neural network StyleFlow choi2024exploiting CVPR24
Audio/Video Transformer LipFD liu2024lips NeurIPS 2024
@@ 3 Dataset Creation
Figure 1: Overview of the TalkingHeadBench dataset creation pipeline. We use portrait images from FFHQ karras2019style as source and audio or video from CelebV-HQ zhu2022celebv as driving signal to generate talking-head deepfake videos using 6 open-source and 1 commercial generators; deepfake images from these detectors are shown at the bottom. We then perform multi-stage data curation to obtain 2984 high-quality deepfake videos along with 2312 real videos. Finally, we design three evaluation protocols to assess detector robustness under distribution shifts in identity and generator characteristics between training and testing sets.
This section details the creation process of our dataset, including data sources, generation methods, cleaning procedures, statistics, and data splits. Fig. 1 provides an overview of this pipeline.
@@ 3.1 Source Data and Generation Overview
To create a diverse deepfake benchmark dataset, we utilize publicly available datasets containing images from FFHQ (Flickr-Faces-HQ) karras2019style and videos of real people and audio of people talking from CelebV-HQ zhu2022celebv , processed by a diverse suite of contemporary deepfake generators. We select six open-source generators: Hallo xu2024hallo , Hallo2 cui2024hallo2 , AniPortrait (Audio-driven) (AniAudio) wei2024aniportrait , AniPortrait (Video-driven) (AniVideo) wei2024aniportrait , LivePortrait guo2024liveportrait , and EMOPortraits drobyshev2024emoportraits , aiming for diversity in both diffusion-based and non-diffusion-based frameworks and driving signal modalities in both video and audio, as summarized in Tab. 2. We choose these generators because they are the latest SoTA generators with publicly available code covering varying modalities and frameworks. Additionally, we create a commercial dataset from MAGI-1 magi1 for experimental testing and show the result in the Supplementary. We also explore other generators such as Hallo3 cui2024hallo3 , X-Portrait xie2024x , and MCNet hong2023implicit , with publicly available code, but do not include them in our dataset since Hallo3 is computationally expensive (GPU memory constraint), and X-Portrait and MCNet generate talking-heads with significant artifacts.

### v2

@@ Abstract
The rapid advancement of talking-head deepfake generation fueled by advanced generative models has elevated the realism of synthetic videos to a level that poses substantial risks in domains such as media, politics, and finance. However, current benchmarks for deepfake talking-head detection fail to reflect this progress, relying on outdated generators and offering limited insight into model robustness and generalization. We introduce TalkingHeadBench, a new benchmark designed to address this gap, featuring talking-head videos from six modern generators, with an additional two emerging generators used exclusively for testing generalization. The dataset is built on an expert-led curation process that filters over 60% of samples to remove videos with noticeable artifacts, presenting a more difficult challenge for detectors. Our evaluation protocols are designed to measure generalization across identity and generator shifts. Benchmarking seven state-of-the-art detectors reveals that models with high accuracy on older datasets like FaceForensics++ show a significant performance drop on our curated data, particularly at strict false positive rates (e.g., TPR@FPR=0.1%). In addition, we identify a trend where detectors focus on background cues instead of facial features using Grad-CAM visualization. Our benchmark aims to accelerate research towards more robust and generalizable detection models in the face of rapidly evolving generative techniques. We release our benchmark and dataset with all data splits and protocols at https://anaxqx.github.io/talkingheadbench.github.io.
@@ 1 Introduction
Figure 1: Overview of the TalkingHeadBench dataset creation pipeline. We use portrait images from FFHQ [karras2019style] as source and audio or video from CelebV-HQ [zhu2022celebv] as driving signal to generate talking-head deepfake videos using 6 open-source generators for our core dataset. The images at the bottom show sample outputs from these six generators as well as from two additional emerging generators (MAGI-1 [magi1] and Hallo3 [cui2024hallo3]) used for evaluating generalization. We then perform multi-stage data curation to obtain 2994 high-quality deepfake videos along with 2312 real videos. Finally, we design three evaluation protocols to assess detector robustness under distribution shifts in identity and generator characteristics between training and testing sets.
@@
Deepfake technology, known for its ability to manipulate facial features and produce highly realistic synthetic videos, has seen rapid adoption in entertainment [liu_deepfacelab_2023, abdolahnejad_deep_2020], marketing [kaur_deepfake_2024], and personalized content creation [qi_mytimemachine_2024, qi_my3dgen_2025, wei_personalized_2025]. Yet, these advancements also introduce serious societal risks. Among deepfake types, talking-head deepfakes, which generate highly realistic head-and-shoulder videos of individuals speaking, are especially concerning. Driven by advanced generative models, these videos are potent tools for spreading misinformation, manipulating public discourse, and executing social engineering attacks [gambin_deepfakes_2024, saini_deepfake_2024, saif_fake_2024]. Their realism and ease of distribution make them particularly dangerous in sensitive domains like politics, finance, and media. In one real-world case, a finance worker was tricked into transferring \$25 million during a video call with a deepfake impersonating the company’s CFO [magramo_finance_2024]. These growing threats highlight the urgent need to understand talking-head deepfakes and to develop robust detection and mitigation strategies [gambin_deepfakes_2024, sirra_csslno_2024, liu_evolving_2025].
@@
In response to the growing threats posed by talking-head deepfakes, numerous facial deepfake benchmarks have been proposed (see Tab. 1), which played a central role in advancing detection methods. However, most existing benchmarks are not well-suited for the modern talking-head landscape. Most deepfake datasets primarily feature face-swapping techniques [rossler2019faceforensics++] rather than talking-head videos, which are generated using different techniques. Although large-scale and collected in the wild, these datasets often contain a high proportion of low-quality samples with noticeable artifacts [chandra2025deepfake, yan2024df40], making it easier to train detectors that ultimately perform poorly on more realistic deepfakes, which pose a greater threat. This leaves a gap in reliably measuring the performance of detectors against the more realistic manipulations produced by the recent talking-head generators. To address these gaps, we introduce TalkingHeadBench—the first manually curated, multi-generator, multi-modal benchmark designed specifically for evaluating detector robustness against state-of-the-art talking-head deepfakes.
@@
Our benchmark features high-quality synthetic videos generated with six academic talking-head generators using both audio- and video-driven diffusion and transformer-based models [drobyshev2024emoportraits, cui2024hallo2, xu2024hallo, wei2024aniportrait, guo2024liveportrait] in the training dataset. To further test generalization, we evaluate models against two emerging generators: an additional academic model (Hallo3 [cui2024hallo3]) and a commercial one (MAGI-1 [magi1]). To ensure dataset quality, we use a five-stage expert curation protocol that removes over 60% of generated videos based on pre-defined artifact criteria. The result is a challenging dataset of 2,994 high-quality talking-head deepfakes from eight academic and commercial generators. This emphasis on quality over quantity creates a more realistic evaluation setting. We also carefully design evaluation protocols to assess the generalization and robustness of state-of-the-art (SOTA) deepfake detectors [dong2023implicit, shao2025deepfake, xu2023tall, xu2024learning, liu2024lips] across identity and generator shifts in training and testing data. In addition to benchmarking, we present a detailed error analysis using Grad-CAM [Selvaraju_2019] to identify common failure cases of SOTA detectors on different generators and inform future improvements in detection strategies. We then note the limitations of SOTA detectors and identify areas of improvement, and provide guidelines on what performance future detectors should aspire to and how to analyze them.
@@
We open-source the TalkingHeadBench benchmark with train-test data for all generators across multiple protocols on our project page. In the near future, we aim to build a more active benchmark repository by a) developing API/code/metrics that also allow authors of new generators to evaluate the ‘detectability’ of their videos; b) setting up two leaderboards comparing generator and detector performance; c) updating the list of generators and detectors every 6 months to accelerate the development cycle.
Our main contributions are:
  • A highly curated benchmark for modern talking-head generation. We release TalkingHeadBench, a multi-modal and multi-generator dataset of 2,994 high-quality deepfake videos, filtered from a much larger pool of deepfake videos using a multi-stage expert-led curation process.
  • An analysis of detector performance at strict thresholds. We show that standard metrics like AUC can obscure poor performance. On our benchmark, SOTA detectors exhibit sharp performance drops at more practical, low-FPR thresholds (e.g., TPR@FPR=0.1%).
  • A study of generalization across identity and generator shifts. Our structured protocols quantify the performance drop of detectors across these distribution shifts, showing the combined effect remains a difficult challenge.
  • An explainability analysis identifying semantic drift. Using Grad-CAM, we show that as generator quality improves, detectors tend to shift their focus from facial regions to background artifacts, exposing a common failure.
In summary, our benchmark demonstrates that while state-of-the-art detectors attain high accuracy on existing deepfake detection benchmarks [rossler2019faceforensics++], their performance drops markedly on TalkingHeadBench. This degradation stems from the inclusion of advanced multi-modal generators, deliberate shifts in identity and generator between training and testing splits, and careful curation that eliminates obvious artifacts in the deepfakes. Together, these factors highlight the need for developing more robust deepfake detectors and establish our benchmark as a foundation for driving future research in this direction.
@@
Table 1: Overview of Facial Deepfake Benchmarks. TalkingHeadBench is the first manually curated, multi-modal, multi-generator benchmark focused on talking-head deepfakes, specifically designed to evaluate detector robustness and generalization. Curation removes $\sim$60% of low-quality samples with obvious artifacts, ensuring a more realistic and challenging benchmark. Generators marked with ✓* use lip-sync-only techniques.
@@
DeepFake-eval-2024 [chandra2025deepfake] N/A Video/Audio/Image 2025 1,208 767
FakeAVCeleb [khalid2021fakeavceleb] Multiple Audio/Video 2021 500 19,500
FaceForensics++ [rossler2019faceforensics++] Multiple Video 2019 1,000 4,000
LAV-DF [cai2022you] ✓* Single Audio/Video 2022 36,431 99,873
AV-Deepfake1M [cai2024av] ✓* Single Audio/Video 2023 286,721 860,039
PolyGlotFake [hou2024polyglotfakenovelmultilingualmultimodal] ✓* Multiple Audio 2024 766 14,472
DF-40  [yan2024df40] Multiple Video/Image 2024 1,590 1M+
TalkingHeadBench (Ours) Multiple Audio/Video 2025 2,312 2,994
@@ 2 Related Work
Deepfake generation has leveraged various facial manipulation techniques over the years. Among them, face swapping methods [nirkin2019fsgan, rosberg2023facedancer, chen2020simswap] are widely used, combining facial regions of various sources to create synthetic videos. They often introduce visible artifacts such as boundary inconsistency, lighting mismatch, or texture irregularity, which modern detectors can learn to exploit easily.
Talking-head generation presents a different set of challenges. Unlike face-swapping, which inserts a swapped facial region and often leaves detectable boundary artifacts, talking-head models synthesize the entire face in motion. These manipulations are driven by audio or video signals and are characterized by their temporal properties and speech-synchronized movements. This makes them a distinct class of forgery that requires evaluation against detectors trained to spot not just spatial inconsistencies, but also subtle temporal and behavioral anomalies. Our benchmark is designed specifically to address this category of deepfake.
@@
Recently, talking-head generation has emerged as a more sophisticated approach. These models synthesize the full face, including expressions and lip motion, from a static portrait using audio or video driving signals by eliminating the obvious seam-based flaws of earlier methods. State-of-the-art diffusion-based talking-head models (see Tab. 2) produce highly realistic outputs with synchronized audio-visual features, significantly raising the detection difficulty.
@@
Despite these advances, existing deepfake benchmarks remain limited in scope. As shown in Tab. 1, most prior datasets either focus on older GAN-based generators or target narrow manipulation types such as face swapping (e.g., FakeAVCeleb [khalid2021fakeavceleb] and FaceForensics++ [rossler2019faceforensics++]) or scrape social media content without controlling of generators (e.g., DeepFake-eval-2024 [chandra2025deepfake]). While newer datasets like LAV-DF [cai2023glitch], AV-Deepfake1M [cai2024av], and PolyGlobFake [hou2024polyglotfakenovelmultilingualmultimodal] incorporate talking-head deepfakes, they mostly emphasize lip-sync-driven synthesis and lack broader modality or generator diversity. Further, these datasets are often massive but uncurated, resulting in many low-quality samples. In contrast, our dataset is built using a curated collection of cutting-edge diffusion-based talking-head generators spanning both audio and video-driven modalities. Through manual inspection, we discarded over 60% of generated videos due to visible artifacts (see Tab. 4), ensuring that only high-quality, realistic samples remain—thereby offering a more meaningful and challenging benchmark.
@@
Recent detection methods (see Tab. 3), though effective on static and unimodal datasets like FaceForensics++ [rossler2019faceforensics++], have not been thoroughly evaluated against multi-modal diffusion-based deepfakes. To fill this gap, we introduce a comprehensive benchmark that reflects the current generation landscape. Our results show that SOTA detectors, which achieve near-perfect accuracy on traditional benchmarks like FaceForensics++, struggle significantly when tested on our dataset, highlighting poor robustness to identity and generator shifts and calling for stronger generalization capabilities.
@@
Table 2: Overview of talking-head generators, highlighted used for creating the train and test sets of TalkingHeadBench.
@@
Image, Audio Diffusion Hallo [xu2024hallo] arXiv24
Image, Audio Diffusion Hallo2 [cui2024hallo2] ICLR25
Image, Audio, Text Diffusion Hallo3 [cui2024hallo3] CVPR25
Image, Video Diffusion X-Portrait [xie2024x] SIGGRAPH24
Image, Video Non-diffusion LivePortrait [guo2024liveportrait] arXiv24
Image, Video Non-diffusion EMOPortraits [drobyshev2024emoportraits] CVPR24
Image, Video Non-diffusion MCNet [hong2023implicit] ICCV23
Image, Video Non-diffusion FSRT [rochow2024fsrt] CVPR24
Image, Audio/Video Diffusion AniPortrait [wei2024aniportrait] arXiv24
Image, Audio/Video Diffusion SkyReels-A1 [qiu2025skyreelsa1expressiveportraitanimation] arXiv25
Image, Text Diffusion MAGI-1 [magi1] N/A
@@
Image CNN SBI [shiohara2022detecting] CVPR22
Image Neural Network CADDM [dong2023implicit] CVPR23
Image Neural Network HiFiNet[hifi_net_xiaoguo]//HiFiNet++[xiao_hifinet_plusplus] CVPR23/IJCV24
Image Transformer DeepFake-Adapter [shao2025deepfake] IJCV25
Image Transformer GenConVit [deressa2025genconvitdeepfakevideodetection] arXiv25
Image Transformer RFFR [shi2025real] PR25
Image Neural network LAA-Net [nguyen2024laa] CVPR24
Video Transformer TALL [xu2023tall]/TALL++ [xu2024learning] ICCV23/IJCV24
Video CNN AltFreezing [wang2023altfreezing] CVPR23
Video CNN/Transformer NACO [zhang2024learning] ECCV24
Video Neural network MM-Det [song2025learningmultimodalforgeryrepresentation] NeurIPS24
Video Neural network StyleFlow [choi2024exploiting] CVPR24
Audio/Video Transformer LipFD [liu2024lips] NeurIPS 2024
@@ 3 Dataset Creation
This section details the creation process of our dataset, including data sources, generation methods, curation procedures, statistics, and data splits. See Fig. 1 for the overview.
@@ 3.1 Source Data and Generation Overview
To create a diverse deepfake benchmark dataset, we utilize publicly available datasets containing images from FFHQ (Flickr-Faces-HQ) [karras2019style] and videos of real people and audio of people talking from CelebV-HQ [zhu2022celebv], processed by a diverse suite of contemporary deepfake generators. We select six open-source generators: Hallo [xu2024hallo], Hallo2 [cui2024hallo2], AniPortrait (Audio-driven) (AniAudio) [wei2024aniportrait], AniPortrait (Video-driven) (AniVideo) [wei2024aniportrait], LivePortrait [guo2024liveportrait], and EMOPortraits [drobyshev2024emoportraits], aiming for diversity in both diffusion-based and non-diffusion-based frameworks and driving signal modalities in both video and audio, as summarized in Tab. 2. We choose these generators because they are the latest SoTA generators with publicly available code covering varying modalities and frameworks.
Additionally, we test generalization against two entirely unseen models: Hallo3 [cui2024hallo3], a recent academic generator, and MAGI-1 [magi1], a commercial text-to-video model. These are used exclusively in our test set to evaluate out-of-distribution robustness. Owing to computational and cost limitations, these test sets are relatively small in size, but they nevertheless provide a meaningful probe of model generalization. We also explore other generators such as X-Portrait [xie2024x] and MCNet [hong2023implicit], but do not include them in our dataset due to significant artifacts observed.
Table 4: TalkingHeadBench dataset statistics across generators.
Data Type # Total # Train # Test Ratio Removed Hallo [xu2024hallo] 420 303 117 $\sim 42\%$ Hallo2 [cui2024hallo2] 432 327 105 $\sim 60\%$ AniPortrait (Audio) [wei2024aniportrait] 542 396 146 $\sim 66.7\%$ AniPortrait (Video) [wei2024aniportrait] 422 314 108 $\sim 80\%$ LivePortrait [guo2024liveportrait] 529 352 177 $\sim 60\%$ EMOPortraits [drobyshev2024emoportraits] 573 381 192 $\sim 50\%$ MAGI-1 [magi1] 66 0 66 $\sim 18\%$ Hallo3 [cui2024hallo3] 10 0 10 $\sim 10\%$ TOTAL Fakes 2994 2073 921 $\sim 63\%$